# GPU-Accelerated DCOPF using Gradient-Based Optimization

### v2

@@ Abstract
DC Optimal Power Flow (DCOPF) is a key operational tool for power system operators, and it is embedded as a subproblem in many challenging optimization problems (e.g., line switching). However, traditional CPU-based solve routines (e.g., simplex) have saturated in speed and are hard to parallelize. This paper focuses on solving DCOPF problems using gradient-based routines on Graphics Processing Units (GPUs), which have massive parallelization capability. To formulate these problems, we pose a Lagrange dual associated with DCOPF (linear and quadratic cost curves), and then we explicitly solve the inner (primal) minimization problem with a dual norm. The resulting dual problem can be efficiently iterated using projected gradient ascent. After solving the dual problem on both CPUs and GPUs to find tight lower bounds, we benchmark against Gurobi and MOSEK, comparing convergence speed and tightness on the IEEE 2000 and 10000 bus systems. We provide reliable and tight lower bounds for these problems with, at best, 5.4x speedup over a conventional solver.
@@ 1 Introduction
Optimal power flow (OPF) is a crucial tool for enabling successful power system operations. It plays a significant role in ensuring the economic operation of power systems by determining the optimal active power output of generators [1]. More specifically, Direct Current (DC) OPF is a simplified version of the OPF problem. It assumes a lossless and unity voltage transmission network, and it can be efficiently solved[2]. DCOPF is the cornerstone for many complex power system optimization problems (e.g., transmission line switching, unit commitment, etc.). A plethora of research has focused on different methods for solving the DCOPF problem [3]. Examples of these methods include barrier-based and Interior-Point Methods (IPMs) [4], commonly used by solvers like Gurobi, MOSEK, and Ipopt, and dual or primal simplex algorithms. While these methods are highly mature, they are serially constrained and do not parallelize effectively [5]. Thus, parallel solving thousands of DCOPF-like problems in a Unit Commitment or Transmission Switching Branch-and-Bound search tree is impractical. While metaheuristics (e.g., Genetic Algorithms, Particle Swarm Optimization) do parallelize, they can also be very slow due to their stochastic nature[6], and they also generally lack convergence guarantees, making them less trustworthy for industry uses.
@@
    Finally, many competitive gradient-based tools have emerged in the literature. We tested and analyzed the sensitivity of three different GBO variants (Adam, AdaGrad, and Gradient with Momentum) for solving our formation, and we identify key hyperparameters and methods that yield swift and reliable convergence on two standard test cases.
@@ 3.2 The Dual Problem of the QP
$\gamma_{q}=\max_{{\tilde{s}}\in K_{*}^{{\rm RSOC}},\mu\geq 0,\lambda}\;-$ $\left\|\hat{c}^{T}+\lambda^{T}\hat{A}+\mu^{T}\hat{C}-\hat{s}^{T}\hat{M}\right\|_{1}$
@@
$\gamma_{q}$ $=\max_{\mu\geq 0,\lambda,s}\quad-\left\|\hat{c}^{T}+\lambda^{T}\hat{A}+\mu^{T}\hat{C}-\hat{s}^{T}\hat{M}\right\|_{1}$
@@
where the vector $s$ is a set of unconstrained dual variables, and $s_{1}=1$ is implied. Numerical testing shows that the objective solutions of (8) and (3.2) exactly agree.
@@ 4 Test Results
In this section, we present the results attained from solving the problem of DCOPF via GBO on CPU and GPU. We compare the performance of the three GBO variants on the LP with regard to iterations and speed. The tests are done on the IEEE 2000 and 10000 bus systems separately [27]. Then, we investigate the effectiveness of using GPUs for solving the two (relatively small and big) problems. At the end, we present an analysis of the hyperparameters associated with the GBO routines and how the tuning process can affect the convergence. All the simulations are run on Apple M2 Pro with 32 GB of shared memory between CPU and GPU, with a total of 12 cores of CPU and 19 cores of GPU and Metal 3 support via Metal.jl [28]. All code is written with minimal memory allocation. Benchmarks were solved with Gurobi and MOSEK 10.
@@ 4.1 Solving the LP with GBO
In the case of the 2k-bus system, AdaGrad and GDM provide a tight solution bound in almost half the best benchmark time (Gurobi), with gaps of 1.09% and almost zero, respectively. These results cannot necessarily be generalized to any system. Therefore, another test is done on the 10k-bus system. For this system, Adam provides a lower bound with a gap of 0.2% in almost one-fifth the time of the best benchmark. Thus, as the size of the system grows, Adam seems to be more effective, while for the smaller system, GDM provides a faster and smoother convergence.
@@
Adam 0.390 1.40 125 0.07
AdaGrad 0.081 0.705 121 1.09
GDM 0.107 0.743 62 4e-4
Gurobi 0.210 - 230 -
@@
Comparing tables 1 and 2, it is worth noting that for the small system, since the solve time is small, the overhead caused by data transfer between CPU and GPU is dominant, so the GPU solve time is slower than the parallelized CPU solve time. However, this is not the case for the large system, and we can see that the iterations run faster on the GPU.
@@ 4.2 Solving the QP with GBO
Looking at the results, GDM gives the fastest and tightest lower bound on the QP problem in the 2k-bus system, while Adam is doing better in the QP problem of the 10k system. The strength of this method is perfectly observable from tables 6 and 2 as we present lower bounds with tightness of 0.2% and 0.44% for the 10k system in half the solve time of the best benchmark.
We benchmark the QP against MOSEK and Gurobi. Tables 5 and 6 show the results for solving the QP on a 2000 and 10000 bus system, respectively. The advantages of using the SOCP dual cone projection are notable. As opposed to the GDM method in LP, using step decay showed promising results for the 10k-bus case for the QP. It enabled faster and more stable convergence in the end phase. The process of tuning the hyper-parameters is very important and greatly affects the convergence of GBO. As can be seen from Fig. 3, for the case of GDM, as we increase momentum, the rise time decreases, but the stability of the convergence is compromised. A balance should be set, considering the priority of the problem solution. Note also how the stability in the first phase and the iterations change for Adam as $\alpha$, $\beta_{1}$ and $\beta_{2}$ change in Fig. 4. The tuned hyperparameters for all the results in this paper are presented in tables 3 and 4.
@@
Table 5: Performance comparison of different algorithms with QP, 2k bus system.
@@
MOSEK 1.628680 28 -

### v3

@@ Abstract
DC Optimal Power Flow (DCOPF) is a key operational tool for power system operators, and it is embedded as a subproblem in many challenging optimization problems (e.g., line switching). However, traditional CPU-based solve routines (e.g., simplex) have saturated in speed and are hard to parallelize. This paper focuses on solving DCOPF problems using gradient-based routines on Graphics Processing Units (GPUs), which have massive parallelization capability. To formulate these problems, we pose a Lagrange dual associated with DCOPF (linear and quadratic cost curves), and then we explicitly solve the inner (primal) minimization problem with a dual norm. The resulting dual problem can be efficiently iterated using projected gradient ascent. After solving the dual problem on both CPUs and GPUs to find tight lower bounds, we benchmark against Gurobi and MOSEK, comparing convergence speed and tightness on the IEEE 2000, 4601, and 10000 bus systems. We provide reliable and tight lower bounds for these problems with, at best, 5.4x speedup over a conventional solver.
@@ 1 Introduction
Optimal power flow (OPF) is a crucial tool for enabling successful power system operations. It plays a significant role in ensuring the economic operation of power systems by determining the optimal active power output of generators [1] [2]. More specifically, DC OPF is a simplified version of the OPF problem. It assumes a lossless and unity voltage transmission network, and it can be efficiently solved[3] [4]. DCOPF is the cornerstone for many complex power system optimization problems (e.g., transmission line switching, unit commitment, etc.). A plethora of research has focused on different methods for solving the DCOPF problem [5] [6]. Examples of these methods include barrier-based and Interior-Point Methods (IPMs) [7], commonly used by solvers like Gurobi, MOSEK, and Ipopt, and dual or primal simplex algorithms. While these methods are highly mature, they are serially constrained and do not parallelize effectively [8]. Thus, parallel solving thousands of DCOPF-like problems in a Unit Commitment or Transmission Switching Branch-and-Bound search tree is impractical. While metaheuristics (e.g., Genetic Algorithms, Particle Swarm Optimization) do parallelize, they can also be very slow due to their stochastic nature[9], and they also generally lack convergence guarantees, making them less trustworthy for industry uses.
@@
    Finally, many competitive gradient-based tools have emerged in the literature. We tested and analyzed the sensitivity of three different GBO variants (Adam, AdaGrad, and Gradient with Momentum) for solving our formation, and we identify key hyperparameters and methods that yield swift and reliable convergence on three standard test cases.
@@ 3.2 The Dual Problem of the QP
$\gamma_{q}=\!\!\!\!\max_{{\tilde{s}}\in K_{*}^{{\rm RSOC}},\mu\geq 0,\lambda}-$ $\left\|\hat{c}^{T}+\lambda^{T}\hat{A}+\mu^{T}\hat{C}-\hat{s}^{T}\sqrt{\hat{M}}\right\|_{1}$
@@
$\gamma_{q}$ $=\max_{\mu\geq 0,\lambda,s}\quad-\left\|{c}^{T}+\lambda^{T}{A}+\mu^{T}{C}-{s}^{T}\sqrt{{M}}\right\|_{1}$
@@
where the vector $s$ is a set of unconstrained dual variables, and all $\hat{\cdot}$ terms have been removed, since $s_{1}=1$ is implied. Numerical testing shows that the objective solutions of (8) and (3.2) exactly agree.
@@ 4 Test Results
In this section, we present the results attained from solving the problem of DCOPF via GBO on CPU and GPU. We compare the performance of the three GBO variants on the LP with regard to iterations and speed. The tests are done on the IEEE 2000, 4601, and 10000 bus systems separately [30]. Then, we investigate the effectiveness of using GPUs for solving the two (relatively small and big) problems. At the end, we present an analysis of the hyperparameters associated with the GBO routines and how the tuning process can affect the convergence. All the simulations are run on Apple M2 Pro with 32 GB of shared memory between CPU and GPU, with a total of 12 cores of CPU and 19 cores of GPU and Metal 3 support via Metal.jl [31]. All code is written with minimal memory allocation. Benchmarks were solved with Gurobi and MOSEK 10.
@@ 4.1 Solving the LP with GBO
In the case of the 2k-bus system, AdaGrad and GDM provide a tight solution bound in almost half the best benchmark time (Gurobi), with gaps of 1.09% and almost zero, respectively. These results cannot necessarily be generalized to any system. Therefore, two other tests are done on the 4601-bus and the 10k-bus system. In the 4601 system, GDM provides a solution of 4.02e-3 gap in almost one-forth of the best benchmark whereas for the 10k system, Adam provides a lower bound with a gap of 0.2% in almost one-fifth the time of the best benchmark. Thus, as the size of the system grows, Adam seems to be more effective, while for the smaller system, GDM provides a faster and smoother convergence.
@@
Adam 0.39 1.40 125 0.07
AdaGrad 0.08 0.70 121 1.09
GDM 0.10 0.74 62 4e-4
Gurobi 0.21 - 23 -
@@
Table 2: Comparison of CPU and GPU LP solver performance for different algorithms (4601-bus).
Algorithm t(s)-CPU t(s)-GPU It Gap(%)
Adam 0.37 0.61 45 0.1
AdaGrad 0.23 0.53 56 6.17e-4
GDM 0.15 0.36 28 4.02e-3
Gurobi 0.63 - 78 -
MOSEK 1.37 - 17 -
@@
Comparing tables 1, 2, and 3, it is worth noting that for the smaller systems, since the solve time is small, the overhead caused by data transfer between CPU and GPU is dominant, so the GPU solve time is slower than the parallelized CPU solve time. However, this is not the case for the large system, and we can observe that the iterations run faster on the GPU as the size of the system grows.
@@ 4.2 Solving the QP with GBO
Looking at the results, GDM gives the fastest and tightest lower bound on the QP problem in the 2k-bus system, while Adam is doing better in the QP problem of the 10k system. The strength of this method is perfectly observable from tables 9 and 3 as we present lower bounds with tightness of 0.2% and 0.44% for the 10k system in half the solve time of the best benchmark.
We benchmark the QP against MOSEK and Gurobi. Tables 7, 8, and 9 show the results for solving the QP on a 2000, 4601, and 10000 bus systems, respectively. The advantages of using the SOCP dual cone projection are notable. As opposed to the GDM method in LP, using step decay showed promising results for the 10k and 4601-bus case for the QP. It enabled faster and more stable convergence in the end phase. The process of tuning the hyper-parameters is very important and greatly affects the convergence of GBO. As can be seen from Fig. 3, for the case of GDM, as we increase momentum, the rise time decreases, but the stability of the convergence is compromised. A balance should be set, considering the priority of the problem solution. Note also how the stability in the first phase and the iterations change for Adam as $\alpha$, $\beta_{1}$ and $\beta_{2}$ change in Fig. 4. The tuned hyperparameters for all the results in this paper are presented in tables 45, and 6.
@@
Table 5: LP and QP Hyperparameter settings for different algorithms (4601-bus).
Algorithm $\alpha$ $\rho$ $\beta_{1}$ $\beta_{2}$
LP Adam 485 - 0.90 0.96
Adagrad 2725 - - -
GDM 6.5 0.94 - -
QP Adam 750 - 0.5 0.999999
AdaGrad 2250 - - -
GDM 2.21 0.8 - -
@@
Table 7: Performance comparison of different algorithms with QP, 2k-bus system.
@@
MOSEK 1.62 28 -
Table 8: Performance comparison of different algorithms with QP, 4601-bus system.
Algorithm t(s)-CPU It Gap(%)
Adam 0.83 156 0.2
AdaGrad 0.34 41 3.8
GDM 0.66 72 4.4
Gurobi 1.24 19 -
MOSEK 1.49 20 -